A telephone company representative estimates that 20% of its customers have call-waiting service. To test this hypothesis, she selected a sample of 100 customers and found that 37% had call waiting. At 5% significance level, is there enough evidence to reject the claim? To give your conclusion,

a. Use the traditional method.

b. Use the p-value method.

c. Use the confidence interval method.
1
Expert's answer
2019-04-30T11:11:43-0400

"a.\;\;Test\;statistic:\;z=\frac{0.37-0.20}{\sqrt{\frac{0.2*(1-0.2)}{100}}}=4.25.\\
Critical\;z-value:\;z_c=1.96.\\
Since\;test\;statistic\;is\;greater\;than\;the\;critical\;value,\;we\;should\;conclude\;that\;\\there\;is\;enough\;evidence\;to\;reject\;the\;claim.\\

b.\;\;P-value\;for\;z=4.25:\;p<0.0001.\\
Since\;P-value\;is\;less\;than\;0.05,\;we\;should\;conclude\;that\;\\there\;is\;enough\;evidence\;to\;reject\;the\;claim.\\
c.\;\;95 \%CI=(0.37-1.96\sqrt{\frac{0.2(1-0.2)}{100}},\;0.37+1.96\sqrt{\frac{0.2(1-0.2)}{100}})=(0.29,\;0.45).\\
Since\;0.2\;does\;not\;lie\;in\;the\;confidence\;interval,\;we\;should\;conclude\;that\;\\there\;is\;enough\;evidence\;to\;reject\;the\;claim.\\"